### v3

# Markovian Scale Prediction: A New Era of Visual Autoregressive Generation

Abstract
00footnotetext: Project Lead, Corresponding Author.
Visual AutoRegressive modeling (VAR) based on next-scale prediction has revitalized autoregressive visual generation. Although its full-context dependency, i.e., modeling all previous scales for next-scale prediction, facilitates more stable and comprehensive representation learning by leveraging complete information flow, the resulting computational inefficiency and substantial overhead severely hinder VAR’s practicality and scalability. This motivates us to develop a new VAR model with better performance and efficiency without full-context dependency. To address this, we reformulate VAR as a non-full-context Markov process, proposing Markov-VAR. It is achieved via Markovian Scale Prediction: we treat each scale as a Markov state and introduce a sliding window that compresses certain previous scales into a compact history vector to compensate for historical information loss owing to non-full-context dependency. Integrating the history vector with the Markov state yields a representative dynamic state that evolves under a Markov process. Extensive experiments demonstrate that Markov-VAR is extremely simple yet highly effective: Compared to VAR on ImageNet, Markov-VAR reduces FID by 10.5% (256×256) and decreases peak memory consumption by 83.8% (1024×1024). We believe that Markov-VAR can serve as a foundation for future research on visual autoregressive generation and other downstream tasks.
Figure 1: Visualization of generated images from our Markov-VAR at 256×256 or 512×512 on ImageNet benchmark.
Figure 2: Observations of the challenges caused by full-context dependency. (a) Comparison of peak computation state (Activations + KV Cache) memory consumption between depth-24 VAR and Markov-VAR on generating 1024×1024 images with a batch size of 25. (b) Metrics and FID performance of VAR under perturbations injected at different scales. MSE, L1 and LPIPS jointly decrease as the perturbation injection scale shifts larger, indicating that early injected perturbations cause greater performance degradation. This is also evidenced by VAR’s largest FID drop at the first injection scale. (c) Residual-Feature Alignment scores (RFA) between each scale and its every previous scale. It is calculated as the cosine similarity between the output residual feature of the current scale and each input feature of all previous scales, combined with $1\times 1$ convolution projection and square root operation, and preserves the directional contribution.
1 Introduction
As a general modeling paradigm, autoregressive modeling has dominated text generation  [66, 67, 61, 8, 1, 36, 52, 51] and multimodal understanding [71, 7, 29]. Yet, due to suboptimal generation quality, it was not widely adopted in visual generation tasks for a long time. Recently, Visual AutoRegressive modeling (VAR) [58] converts next-token prediction to next-scale prediction, further unleashing the potential of the autoregressive modeling to generate high-quality visual content, making autoregression once again a mainstream modeling paradigm for visual generation [26, 49, 44, 18, 22, 80]. However, VAR’s full-context dependency, requiring attention over all previous scales for next-scale prediction, can introduce certain challenges:
I. Substantial computational cost. As the scale increases, the number of tokens grows quadratically. Moreover, cumulative modeling across multiple previous scales in VAR accelerates the superlinear increase in computational cost, as shown in Figure 2 (a). Such substantial computational cost not only slows down training and inference, but also severely limits VAR’s scalability and practicality.
II. Continuous error accumulation. As a chain-based modeling paradigm, autoregression’s unidirectional causality chain prevents early prediction errors from being corrected and leads to their continuous propagation. Figure 2 (b) shows that early perturbations impact the performance more severely than late ones, indicating that errors continuously accumulate during error propagation. In addition, VAR exacerbates the error accumulation in two aspects. First, its full-context dependency repeatedly utilizes and iteratively accumulates errors from previous scales. Second, rapid growth of the token sequence length extends the chain of error propagation and accumulation. This cross-scale error accumulation significantly degrades both the quality and stability, especially for high-resolution visual generation.
III. Cross-scale interference. VAR’s coarse-to-fine modeling requires learning distinctive scale-specific representations at each scale. However, for the current scale, full-context dependency leads attention to aggregate all previous scales, where mixed information of multiple scales makes attention and gradients from different scales compete or conflict with each other in the shared feature space, thereby suppressing the learning of distinctive representations at the current scale and limiting the improvement of generation quality. As illustrated in Figure 2 (c), we compute RFA scores to quantify the directional contribution of the input feature from previous scales to the current predicted residual feature, further measuring cross-scale consistency in learning the current representation. We find that early scales typically have a negative impact on learning distinctive representations at the current scale.
These challenges motivate us to explore whether it is feasible to develop a visual autoregressive generation model without full-context dependency that outperforms traditional VAR in both performance and efficiency. Inspired by the concept of sufficient statistics [59, 60, 2] in information theory, we suppose that each node in continuous chain-based propagation inherently maintains historical information to a certain extent. Appropriately distilling this into a representative dynamic state can achieve effective prediction without relying on all original historical information.
In light of this motivation, we develop Markov-VAR, a visual autoregressive generation model with Markov process. Markov-VAR refines next-scale prediction into Markovian scale prediction, where each scale prediction depends only on the current one rather than all previous scales. Chain-based unidirectional autoregressive modeling makes the current scale already encoding representative historical information for feature prediction [69, 73, 37]. Thus, we can treat each scale as a Markov state. However, non-full-context dependency inevitably discards substantial original historical information compared to full-context dependency. We establish a lightweight history compensation mechanism to enrich the Markov state for historical information. Specifically, we employ a sliding window that compresses certain previous scales within it into a compact history vector, which is then integrated with the Markov state to construct a representative dynamic state. By modeling these dynamic states as a Markov process, Markov-VAR achieves Markovian scale prediction.
Extensive experiments demonstrate Markov-VAR’s superior performance and efficiency. For image generation on ImageNet [10], compared to VAR, Markov-VAR reduces FID from 3.61 to 3.23 (256×256) and remarkably decreases peak memory consumption of the computation state from 117.9GB to 19.1GB (1024×1024). As a foundation model, its performance and efficiency may become more promising when combined with other enhancement or acceleration techniques. Moreover, Markov-VAR outperforms many alternative-paradigm models or improved VAR-like variants in various comparisons. Our contributions are as follows:
  • We advance visual autoregressive generation by exploring modeling without full-context dependency, refining next-scale prediction as Markovian scale prediction to reformulate VAR as a non-full-context Markov process.
  • We propose Markov-VAR with a history compensation mechanism to mitigate the historical information loss in the Markov process. Markov-VAR comprehensively surpasses VAR in both performance and efficiency, extremely simple yet highly effective.
  • We publicly release the full series of Markov-VAR model weights, hoping to serve as a foundation model and facilitate future research on various tasks.
2 Related Work
Visual Autoregressive Generation.
Early visual autoregressive generation models such as PixelRNN [63], PixelCNN [62] and PixelCNN++ [46] achieved pixel-level conditional modeling, yet this paradigm suffers from limited long dependency and extremely slow sampling. Subsequent works [43, 13, 64, 82, 12, 28, 76, 54, 65] treating image synthesis as a token prediction task relied on Vector Quantization (VQ) [64] to discretize continuous feature maps and autoregressively modeling visual token sequence. However, their raster-scan prediction order conflicts with the spatial structure of the images, leading to inferior performance compared to other alternative-paradigm models [25, 17, 53, 19, 20, 50, 15, 5]. Recently, VAR [58] reformulates next-token prediction as next-scale prediction, further improving the generation quality of visual content. Some VAR-like variants, such as Infinity [18], HART [57] and VAR-CLIP [78] extend VAR to text-to-image generation. Other further improved VAR-like models [32, 16, 42, 72, 70, 23, 34, 83] perform various tasks, such as image super-resolution [41], 3D Object Generation [6] and image editing [68].
Chain-based Markov Modeling.
The Markov assumption [38] that posits that the current state depends only on its immediate predecessor, has been widely employed for research in deep learning. Diffusion-based models [9, 79, 81, 39] have achieved remarkable success in various tasks based on this principle. Chain-based Markov modeling has been introduced to alleviate efficiency challenges of long-context modeling. MARCOS [33] and MCoT [74] introduce structured memory nodes that reformulate chain-of-thought reasoning into a Markovian process for more lightweight inference. In visual autoregressive modeling, although HMAR [26] pioneeringly introduces Markov dependency to improve generation performance, it comes at the cost of increasing inference steps and the token sequence length. MVAR [77] innovatively introduces spatial-Markov attention, focusing on reduce computational complexity. We argue that historical information preservation, more concise modeling, and more comprehensive scaling validation are essential for further advancing Markov visual autoregressive modeling. Therefore, we revisit these challenging trade-off problems and employ chain-based Markov modeling to develop a visual autoregressive foundation model that reconciles performance with efficiency.
3 Method
3.1 Preliminary: Next-Scale Prediction
VAR [58] reformulates autoregressive modeling from next-token prediction to next-scale residual feature prediction, which generates images in a coarse-to-fine manner based on full-context dependency. The modeling process involves $T$ multi-scale residual features $\{R_{1},R_{2},\cdots,R_{T}\}$ defined by a size set $\{S_{1}\times S_{1},S_{2}\times S_{2},\cdots,S_{T}\times S_{T}\}$. At the $t$-th scale, VAR predicts the residual feature $R_{t}\in\mathbb{R}^{S_{t}\times S_{t}}$ based on all previous residual scales $R_{<t}=\{R_{1},R_{2},\cdots,R_{t-1}\}$. The autoregressive likelihood is formulated as follow:
$p(R_{1},R_{2},\cdots,R_{T})~=~\prod_{t=1}^{T}p(R_{t}~\bigl|~\langle\mathrm{sos}\rangle,R_{<t}),$ (1)
where $R_{<t}$ denotes the “prefix” of $R_{t}$, and $\langle\mathrm{sos}\rangle$ is the start conditional embedding.
Figure 3: Left: Comparison of modeling process between VAR and Markov-VAR when predicting the 6-th scale & Comparison of visual context between next-scale prediction and Markovian scale prediction during generation. Markov-VAR utilizes a history compensation mechanism to enrich the current scale for historical information. Right: The overall framework of Markovian scale prediction with Markov-VAR Transformer. [S] is the start token with condition embedding.
3.2 Markovian Scale Prediction
In contrast to VAR, Markov-VAR is based on Markovian scale prediction. The prediction process treats each scale as a Markov state and models it as a chain-based Markov process. The size set $\{S_{1}\times S_{1},S_{2}\times S_{2},\cdots,S_{T}\times S_{T}\}$ defines $T$ multi-scale residual features $\{R_{1},R_{2},\cdots,R_{T}\}$. At the $t$-th scale, Markov-VAR predicts the Markov state $M_{t}\in\mathbb{R}^{S_{t}\times S_{t}}$ based on the Markov state $M_{t-1}$ (the current residual feature $R_{t-1}$). Therefore, Markovian autoregressive likelihood can be formulated as follows:
$p(R_{1},R_{2},\cdots,R_{T})~=~\prod_{t=1}^{T}p(R_{t}~\bigl|M_{t-1}),$ (2)
where Markovian state $M_{t}=f_{\phi}(R_{t},M_{t-1})$ is a representative dynamic state for the residual feature $R_{t}$, and $f_{\phi}(~)$ is a state update function, $M_{0}=\langle\mathrm{sos}\rangle$.
3.3 Markov-VAR
Figure 3 illustrates the comparison between VAR and Markov-VAR, and more details of Markov-VAR.
Markov State.
Classical information theory [59] and its applications in deep learning [60, 2] indicate that: information $I(c_{<t};c_{t})$ of the complete history $c_{<t}$ with the current time $c_{t}$ is highly redundant. There exists a sufficient statistic $c_{t-1}$ such that: $I(c_{t-1};c_{t})=I(c_{<t};c_{t})$. Therefore, we argue that continuous chain-based modeling can propagate this sufficient statistic, enabling effective prediction of subsequent scales only based on the current scale. This theory allows us to naturally treat each scale as a Markov state.
History Compensation Mechanism.
While each scale can be directly treated as a Markov state and modeled in a Markov process, it must be acknowledged that Markovian scale prediction performs suboptimally compared to next-scale prediction based on full-context dependency when the current scale cannot possess the complete sufficient statistic or substantial historical information is discarded. To maintain the Markov process while further improving prediction performance, we design a history compensation mechanism to supplement historical information for prediction.
We set a sliding window $\mathcal{W}$ of size $N$ to store the previous $N$ continuous scales. For the $t$-th residual scale $R_{t}$, the corresponding sliding window $\mathcal{W}_{t}$ is denoted as:
$\mathcal{W}_{t}=\{E_{t-1},E_{t-2},\dots,E_{t-N}\},$ (3)
where $E_{t}\in\mathbb{R}^{n^{t}\times d}$ denotes the embedded feature at the $t$-th scale, which is obtained through word embedding and up-interpolation of the residual scale $R_{t-1}$. $n^{t}$ is the number of tokens at the $t$-th scale and $d$ is the dimension of token.
We define the token sequence of $E_{t}$ as $X_{t}\in\mathbb{R}^{n^{t}\times d}$. Thus, the overall token sequence $\hat{X}_{t}$ of all scales within the sliding window is obtained by a concat operation: $\hat{X}_{t}=Concat(X_{t-1},X_{t-2},\dots,X_{t-N})$. We aggregate $\hat{X}_{t}$ into a fixed-dimensional history vector $h_{t}$ via cross-attention:
$h_{t-1}=\text{Attn}(q,\hat{X}_{t},\hat{X}_{t}),$ (4)
where $q$ is a learnable query representing the global state, and $\hat{X}_{t}$ provides the key–value pairs.
When predicting the $t$-th residual scale $R_{t}$, we broadcast the history vector $h_{t-1}$: $H_{t-1}=\mathbf{1}_{n_{t-1}}h_{t-1}^{\mathsf{T}}$, and concatenate history vectors $H_{t-1}$ and feature scale $E_{t-1}$ element-wise to obtain a representative dynamic state $M_{t-1}$:
$M_{t-1}=Concat(E_{t-1},H_{t-1}).$ (5)
After modeling the evolution of representative dynamic states as a Markov process, Markov-VAR starts to operate.
3.4 Training Strategy
Following the conventional autoregressive paradigm, we construct the ground-truth residual feature sequence $R$ as the supervision target for Markovian scale prediction.
Markov-VAR is trained under the teacher-forcing scheme while further adopting the Markovian attention shown in Figure 3, restricting each scale to attend only to its current state. Other main strategies remain consistent with standard visual autoregressive training.
Algorithm 1 shows the training process, where modules with subscript $\theta$ are learnable. $\hat{f}$ denotes the final target feature of Markov-VAR. $\mathcal{F}=[\langle\mathrm{sos}\rangle,X_{1},\cdots,X_{T-1}]$ represents a queue of token sequences. The overall scale sequence $\hat{R}$ consist of the residual features predicted for each scale. Finally, the training loss $\mathcal{L}$ is calculated by the cross-entropy (CE) [48] between the ground-truth $R$ with $\hat{R}$.
Input: $R_{1},\dots,R_{T}]$, NN, $(S_{t}\times S_{t})_{t=1}^{T}$
Output: Trained model parameters $\theta$
$M=[M_{0},\cdots,M_{T-1}]$,    $M_{0}\leftarrow\langle\mathrm{sos}\rangle$ ;
$\mathcal{F}$ $\leftarrow$ Queue();    $\mathcal{W}$ $\leftarrow$ Queue();
$Queue\_Push(\textnormal{{$\mathcal{F}$}},\langle\mathrm{sos}\rangle)$;    $\hat{f}\leftarrow tensor(0)$ ;
forall $t=1$ to T1T-1 do
 $\hat{f}=\hat{f}+Up(R_{t},S_{T})$ ;
 // $\triangleright$ Add residual
 $E_{t}=Embed_{\theta}(Down(\hat{f},S_{t+1}))$ ;
 $Queue\_Push(\textnormal{{$\mathcal{F}$}},E_{t})$
end forall
forall $t=1$ to T1T-1 do
 if $\textnormal{{$\mathcal{W}$}}.size()==N$ then
    $Queue\_Pop(\textnormal{{$\mathcal{W}$}})$ ;
    // $\triangleright$ Slide window
   end if
 $X_{t}=Queue\_Front(\textnormal{{$\mathcal{F}$}})$ ;
 $Queue\_Pop(\textnormal{{$\mathcal{F}$}})$ ;
 $Queue\_Push(\textnormal{{$\mathcal{W}$}},X_{t})$ ;
 // $\triangleright$ Slide window
 $\hat{X}_{t}=Concat(\textnormal{{$\mathcal{W}$}})$ ;
 $H_{t}=Broadcast(Attn(q_{\theta},\hat{X}_{t},\hat{X}_{t}))$ ;
 $M_{t}=Concat(\textnormal{{$\mathcal{F}$}}.Front(),H_{t})$ ;
 // $\triangleright$ Update
end forall
$\hat{R}_{1},\cdots,\hat{R}_{t}]\leftarrow Model_{\theta}([M_{0},\cdots,M_{T-1}])$ ;
$\mathcal{L}=\sum_{t=1}^{T}CE(\hat{R}_{t},R_{t})$ ;
// $\triangleright$ Compute loss
$\mathcal{L}.Backward()$ ;
// $\triangleright$ Update $\theta$ via $\mathcal{L}$
return $\theta$
Algorithm 1 Training Process of Markov-VAR
4 Experiment
4.1 Experiment Settings
Models and Baselines.
We construct Markov-VAR models with depths $d$ of 16, 20, and 24 layers, the network structure is set as: the width $w=64d$, the number of attention heads $h=d$ and the dropout rate $dr=0.1\cdot d/24$. We utilize Rotary Positional Embedding [55] for learnable positional encoding, and adopt LLaMA-style [61] attention and MLP blocks following [57]. For image tokenization, we employ the pre-trained multi-scale VQ-VAE tokenizer from VAR. To evaluate the effectiveness of our design, we compare Marokov-VAR with the VAR [58] and other VAR-like models on comparable model sizes. We further include comparisons with broad generative models of alternative paradigms, such as diffusion models, GANs, and next-token prediction autoregressive models, to provide a comprehensive evaluation.
Datasets and Metrics.
We conduct the pre-training of Markov-VAR on the ImageNet-1K dataset [10], following the standard class-to-image generation model setting. For quantitative evaluation, we use Fréchet Inception Distance (FID), Inception Score (IS), Precision, and Recall as the primary metrics to comprehensively assess generation quality. We evaluate efficiency in terms of inference time and peak GPU memory consumption of computation state, which is mainly composed of Transformer activations and KV cache.
Implementation Details.
All models are trained in similar settings, with a base learning rate of $8\times 10^{-5}$, using the AdamW optimizer [35] with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. The batch size ranges from 768 to 1536, and the training epochs vary from 200 to 400, depending on the model depth. Training is implemented on 8 NVIDIA H200 GPUs. Evaluation is conducted on a single NVIDIA H200 GPU.
Table 1: Performance comparison on ImageNet 256×256 class-conditional generation with VAR and VAR-like models. Evaluation metrics include Fréchet Inception Distance (FID) and Inception Score (IS). Precision and Recall jointly assess the fidelity–diversity trade-off of generated images. $\downarrow$’ and $\uparrow$’ indicate that lower or higher values are preferable.
[…]
M-VAR-d20 [44] 900M 2.41 308.4 0.85 0.58
FlexVAR-d24 [23] 1.0B 2.21 299.1 0.83 0.59
HMAR-d24 [26] 1.3B 2.10 324.3 0.83 0.60
NestAR-H [70] 1.3B 2.22 342.4 0.79 0.57
VAR-d16[58] 310M 3.61 225.6 0.81 0.52
VAR-d20[58] 600M 2.67 254.4 0.81 0.57
VAR-d24[58] 1.0B 2.17 271.9 0.81 0.59
VAR-d30[58] 2.0B 2.14 275.4 0.80 0.60
[…]
Table 2: Board performance comparison on class-to-image generation on ImageNet $256\times 256$ benchmark with more alternative-paradigm models. Step metric accounts for the number of sampling steps during inference. The suffix ‘-re’ denotes rejection sampling.
[…]
GAN BigGAN [3] 112M 6.95 224.5 0.89 0.38 1
GAN GigaGAN [24] 569M 3.45 225.5 0.84 0.61 1
GAN StyleGAN-XL [47] 166M 2.30 265.1 0.78 0.53 1
Diffusion ADM [11] 554M 10.94 101.0 0.69 0.63 250
Diffusion CDM [21] - 4.88 158.7 - - 8100
Diffusion LDM-4 [45] 400M 3.60 247.7 - - 250
Diffusion DiT-XL/2 [40] 675M 2.27 278.2 0.83 0.57 250
Masked AR MaskGIT [4] 227M 6.18 182.1 0.80 0.51 8
Masked AR MaskGIT-re [30] 227M 4.02 355.6 - 8
Masked AR MAGE [31] 230M 6.93 195.8 - - 20
Next-token AR VQGAN [14] 227M 18.65 80.4 0.78 0.26 256
Next-token AR VQGAN  [14] 1.4B 15.76 74.3 - - 256
Next-token AR VQGAN-re [75] 1.4B 5.20 280.3 - - 256
Next-token AR ViT-VQGAN [75] 1.7B 4.17 175.1 - - 1024
Next-token AR ViT-VQGAN-re [75] 1.7B 3.04 227.4 - - 1024
Next-token AR RQTran [27] 3.8B 7.55 80.4 0.78 0.26 68
Next-token AR RQTran-re [27] 3.8B 3.80 323.7 - - 68
Next-token AR LlamaGen-B [56]  [^1] M 5.46 193.6 0.83 0.45 256
Next-token AR LlamaGen-L [56] 343M 3.81 248.3 0.83 0.52 256
Next-token AR LlamaGen-XL [56] 775M 3.39 227.1 0.81 0.54 256
Next-token AR LlamaGen-XXL [56] 1.4B 3.09 253.6 0.83 0.53 256
[…]
4.2 Main Results
Generation Performance.
We first evaluate the generation performance of Markov-VAR. We conduct comparsion with VAR and other VAR-like models [18, 44, 26, 70] on ImageNet $256\times 256$ class-conditional image generation. As shown in Table 1, Markov-VAR demonstrates a comprehensively better performance, achieving higher generation quality than other VAR models of comparable size. For example, on the depth-16 model, Markov-VAR improve the FID performance from 3.61 to 3.23 (10.5%) and IS performance from 225.6 to 256.2 (13.6%) compared to VAR. Even on Precision and Recall metrics, Markov-VAR matches or surpasses VAR. Markov-VAR’s advantages can be observed across other sizes. Compared with M-VAR-d20, Markov-VAR-d20 achieves compelling performance while using only 70% of its parameters, demonstrating that Markov-VAR is both effective and parameter-efficient.
To further demonstrate the performance advancements of Markov-VAR. We conduct a board comparison on class-to-image generation on ImageNet $256\times 256$ benchmark with more alternative-paradigm models, including GANs [3, 24, 47], diffusion models [11, 20, 45, 40], mask autoregressive models [4, 30, 31], and next-token prediction autoregressive models [56, 14, 75, 27]. As shown in Table 2, Compared with Diffusion models, Markov-VAR achieves better overall performance in terms of both parameter efficiency and generation quality. Compared to the size-similar next-token autoregressive model LlamaGen-L, Markov-VAR achieves superior performance across all evaluation metrics. In contrast to other autoregressive models, Markov-VAR maintains competitive generation quality while requiring fewer sampling steps, demonstrating its higher inference efficiency. Further comparisons in the table also reveal Markov-VAR’s favorable generation performance and parameter efficiency.
Generation Efficiency.
[…]
Table 3: Comparison of inference time and peak memory between Markov-VAR and other VAR-like models across various resolutions and depths. Peak GPU memory is reported as the average of 5 runs with a batch size of 25 on a single NVIDIA H200. Inference time is measured in seconds per batch on a NVIDIA H200.
Model Depth Res. Time(s) $\downarrow$ Memory(GB) $\downarrow$
VAR [58] 16 256 0.303 5.8
VAR [58] 16 512 0.824 18.4
VAR [58] 16 1024 3.303 66.7
HMAR [26] 16 256 0.309 3.1
HMAR [26] 16 512 0.909 5.6
FlexVAR [23] 16 256 0.395 6.8
FlexVAR [23] 16 512 0.944 20.1
FlexVAR [23] 16 1024 3.409 77.9
[…]
VAR [58] 24 256 0.711 12.4
VAR [58] 24 512 1.335 31.4
VAR [58] 24 1024 5.891 117.9
[…]
4.3 Ablation and Analysis
Ablation on History Compensation Mechanism.
To verify the effectiveness of our sliding-window-based history compensation mechanism, we conduct ablation and further compare ours to other mechanisms. Global history is a full-context compensation, which continuously fuses and updates all previous scales, and hybrid history combines both full-context and non-full-context compensation. In Table 4, all types of compensation mechanism improve generation quality compared to the one without history compensation. However, our sliding-window-based history compensation mechanism achieves the overall best preservation of historical information. These validate the effectiveness of our design and are consistent with the analysis in Section 3.3.
Table 4: History compensation ablation on the depth-16 model across different history compensation mechanisms.
Method Depth Param FID $\downarrow$ IS $\uparrow$
w/o History 16 300M 3.64 247.7
Global History 16 324M 3.41 245.2
Hybrid History 16 359M 3.45 257.4
[…]
Ablation on Sliding Window Size.
To determine a suitable sliding window size for Markov-VAR’s history compensation mechanism, we conduct an ablation study on Markov-VAR with depths of 16 and 20. As shown in Table 5, models with the two depths consistently achieve their best performance when window size is set to 3. In comparison, when the window size is set to 1, Markov-VAR-d16 respectively decreases FID and IS to 3.53 and 237.8. All these results are also consistent with the analysis of cross-scale interference discussed in Introduction that full-context dependency may hinders distinctive scale feature learning, and the most recent three scales typically have a positive effect on learning.
Table 5: Ablation study of the sliding-window-based history compensation mechanism with different window sizes.
Window Size FID(d16) $\downarrow$ IS(d16) $\uparrow$ FID(d20) $\downarrow$ IS(d20) $\uparrow$
1 3.53 237.8 2.50 267.9
2 3.39 248.6 2.47 281.4
3 3.23 256.2 2.44 286.1
4 3.33 252.3 2.56 278.2
Figure 4: Analysis of peak memory consumption of Markov-VAR across various depths and resolutions at different scales.
Memory Consumption Analysis.
As shown in Figure 4, we investigated the peak memory consumption of computation state across various depths and resolutions at different scales. Compared with VAR, Markov-VAR exhibits a less steep scaling trend rather than the exponential growth trend observed in VAR. As the model size and target resolution increase, the efficiency advantage of Markov-VAR becomes increasingly evident. Most importantly, because Markov-VAR follows a Markovian modeling process, it does not require any KV-cache computation, which fundamentally accounts for its significantly lower computational cost. Moreover, even the computation of Transformer activation in VAR is also larger than that of Markov-VAR, further demonstrating the computational efficiency advantage of Markov-VAR. All these results demonstrate that Markovian Scale Prediction effectively alleviates the exponential growth of full-context modeling, making it more promising for scaling and generating visual content at larger scales.
Figure 5: Scaling law analysis of Markov-VAR between performance metrics and model sizes with power-law fitted equations.
Figure 6: Visualization of generation process in Markov-VAR at $256\times 256$ resolution.
Analysis of Scaling Law.
We vary the depth of Markov-VAR from 6 to 24 to analyze the scaling law, corresponding to parameter scales ranging from 19.80M to 1.02B. In Figure 5, charts (a) to (d) show the fitted scaling relation between model size (log10) and performance metrics. Both loss and error rate consistently decrease as model size increases, following obvious power-law trends with high coefficients of determination ($R^{2}>0.99$), indicating that Markov-VAR follows the expected scaling law.
4.4 Visualization
Main Generation Visualization.
Figure 1 shows representative image samples generated by our Markov-VAR. Given the scale and quality of ImageNet, Markov-VAR performs reasonably well. Its generation quality may further improve when trained on larger and higher-quality datasets.
Visualization of Generation Process
We visualize the Markovian scale prediction process in Figure 6. This visualization provides an intuitive understanding of how Markov-VAR constructs a semantic structure and fills in fine-grained details across multiple scales. The earliest scales capture only very coarse global structures. As the scale increases, the model gradually refines the semantics. Later scales enrich high-frequency textures and refine local details such as fur patterns and cloud shapes. Notably, the refinement is both smooth and consistent, demonstrating that each Markovian scale effectively preserves essential historical information without relying on full-context dependency.
Visual Comparison between VAR and Markov-VAR.
Figure 7 shows that both models perform well on certain generated samples, even in some cases, Markov-VAR even produces better semantics and higher quality than VAR.
Figure 7: Visual comparison between VAR and Markov-VAR.
5 Conclusion
In this paper, we address the challenges of VAR’s full-context dependency by reformulating it as a Markov process. We propose Markov-VAR, which employs Markovian scale prediction by treating each scale as a Markov state and modeling its evolution as a Markov process. Experiments and analysis demonstrate that Markov-VAR achieves effective and efficient visual generation.